Angle A and angle B are supplementary angles and angle A is 15 degrees less than four times angle B. Find the 2 angles.
Answer by Earlsdon(6098)

If A and B are suplementary angles, then:
A+B = 180 and, from the problem:
A = 4B-15 Substitute this into the above equation and solve for B.
(4B-15)+B = 180 Simplify.
5B-15 = 180 Add 15 to both sides.
5B = 195 Divide both sides by 5.
B = 39 degrees
A = 180-39 Supplementary angles.
A = 141 degrees.